OpenStudy (ubie01):

Factor completely. -2k - k3 - 3k2 k(-k + 1)(k - 2) -k(k - 1)(k - 2) -k(k + 1)(k + 2)

OpenStudy (ubie01):

@hartnn

hartnn (hartnn):

first factor out the 'k' from all the terms, what do you get?

OpenStudy (ubie01):

I'm not sure. I'm not very good with this kind of math.

hartnn (hartnn):

\(\large ☻☺+☻♥= ☻(☺+♥) \\ \text{this is factoring, } \\ \text{☻ was common from both terms and factored out.}\) So in, \(-k^3 -3k^2 -2k\) there is one 'k' which is common in all the terms! factor that out.

OpenStudy (ubie01):

-k(k^2 + 3k - 2)?

hartnn (hartnn):

very good try! from -2k, if we factor -k, we get +2, so it is \(\Large -k(k^2 +3k+2)\)

hartnn (hartnn):

now, can you tell me 2 numbers whose sum is +3 and product is +2 ?

hartnn (hartnn):

what are you confused about?

OpenStudy (ubie01):

I'm not good with polynomials. I haven't been good since I've started studying them.

OpenStudy (skullpatrol):

It takes a lot of practice...

OpenStudy (skullpatrol):

...have you tried the khan academy?

OpenStudy (skullpatrol):

go back to the very beginning definition of a polynomial and work your way through the videos

hartnn (hartnn):

by that time, go through this steps and let me know if any doubt: \(k^2 +3k +2 = k^2 + 2k+k+2 = k(k+2) + 1(k+2) = (k+1)(k+2)\) so yes, -k(k+1)(k+2) is indeed correct.